OpenStudy (anonymous):

factor by grouping, Assume any variable exponents represent whole numbers x^3-4x^2+5x-20

OpenStudy (anonymous):

Is there a common factor you see in \(x^3-4x^2\)? What about in \(5x-20\)?

OpenStudy (anonymous):

the x and that 20 can be divided into 5

OpenStudy (anonymous):

which ever way that goes

OpenStudy (anonymous):

@criddle3583 only one \(x\) in common in \(x^3-4x^2\)? for the \(5\), you're right -- \(5x-20=5(x-4)\). We're getting close.

OpenStudy (anonymous):

and x+5 so its (x-4)(x+5)

OpenStudy (anonymous):

Nope @criddle3583... not quite... :-)

OpenStudy (anonymous):

We can actually find that \(x^2\) is in common in \(x^3-4x^2\), so we can factor it out to yield \(x^3-4x^2=x^2(x-4)\). So our entire expression is then just$$x^2(x-4)+5(x-4)$$-- hey, notice we have an \(x-4\) in common which we can factor it. Do so:$$(x-4)(x^2+5)$$

OpenStudy (anonymous):

ok thank you :)